A tuning fork generates sound waves with a frequency of 232 Hz. The waves travel in opposite directions along a hallway, are reflected by walls, and return. The hallway is 42.0 m long and the tuning fork is located 14.0 m from one end. What is the phase difference between the reflected waves when they meet at the tuning fork? The speed of sound in air is 343 m/s.
Answer #1

१५+ 28 = 42m 28 m 140 = २8m → - -- x2= som A-- = ---- = S6 - 28 = 28 m v_ 343 m}s - 1.4784m 232 Hz Alk 25 - = 18.938 .4784 1.